This circle is centered at the origin, and the length of its radius is 8. What is the equation of the circle?
i know the equation is x(exponent)2+y (exponent)2= r (exponent)2
r stands for radius but i dont know if i have to do 8X8 ... or just leave it like that when solving the equation

Answer by checkley77(12569)

x^2+y^2=8^2
x^2+y^2=64